OpenStudy (vitalina):

Consider a NUMA system with local memory latency=1τ and remote memory latency=4τ. If a program execution results in 80% local access and 20% remote access, what is the average memory latency?

OpenStudy (rsmith6559):

( 8 accesses * 1 ) + ( 2 accesses * 4 ) = 16 / 10 accesses = 1.6 would be my guess.

OpenStudy (vitalina):

thanks. that's what I was thinking but wasn't sure :)